OpenStudy (anonymous):

For 2x - 3y = -7, when x = 4, What is the value of y?

OpenStudy (anonymous):

2(4) - 3y = -7 -3y = -7 - 8 -3y = 15 -3y/3 = 15/-3 y = -5

OpenStudy (anonymous):

2x - 3y = -7 2(4) - 3y = -7 8 - 3y = -7 -3y = 1 y = -3

OpenStudy (anonymous):

@crackdope You have to ADD 8 to both sides not SUBTRACT because 8 is POSITIVE NOT NEGATIVE.

OpenStudy (anonymous):

the answer is 5 guys

OpenStudy (anonymous):

and becuase 8 is positive you would subtract it from both sides to get rid of it, so that is why it is SUBTRACT, not add

OpenStudy (anonymous):

2x- 3y= -7 2(4) - 3y= -7 8- 3y= -7 -8 -8 ------------- -3y= -7-8 -3y= - 15 ---- ----- -3 -3 y= -15/-3 y=5

OpenStudy (anonymous):

preciate it